OpenStudy (anonymous):

hey!How do you find the equation of the tangent line (x-1)^2=6(y+2) @(-5,4)....the derivative of the given eqn is 1/3x-1/3 btw:)

OpenStudy (amistre64):

you wanna derive this?

OpenStudy (amistre64):

(x-1)^2=6(y+2)

OpenStudy (anonymous):

i already did that..n the deriv is the slope right..jus wonderin wat u do with the given point to find the equation at that point

OpenStudy (amistre64):

if the derivative is what you posted; then input (x=-5) and solve to get the slope for your line

OpenStudy (amistre64):

y-4 = f'(-5) (x+5)

OpenStudy (anonymous):

ohh...got it! i got y-4=-2(x+5):)

OpenStudy (amistre64):

that looks to be right :) good job

OpenStudy (anonymous):

thank u so much again!:) u rock